OpenStudy (anonymous):

If g(x) = x^2 + 3, find g(4).

OpenStudy (anonymous):

g(x) = x^2 + 3 g(4) = 4^2 + 3 what do you get

OpenStudy (anonymous):

19?

OpenStudy (anonymous):

16 + 3 = 19 correct